OpenStudy (anonymous):

PLEASE HELP WILL FAN & MEDAL?! which logarithm is equivalent to log3^10 - log3^2 a)log320 b)log38 c)log35 d)log312

OpenStudy (anonymous):

the threes are all little btw

OpenStudy (anonymous):

@welshfella ill medal again for your help?

OpenStudy (aveline):

To subtract two logarithms, you divide them: logb(m) – logb(n) = logb(m/n) log3(10)-log3(2)=log3(10/2) Can you simplify?

OpenStudy (anonymous):

log35, so option c??

OpenStudy (aveline):

Correct

OpenStudy (anonymous):

thanks i appreciate it